OpenStudy (anonymous):

How do you solve x^2-13x+36

OpenStudy (anonymous):

So far I have x^2=x 36 is 62ab=2(x)(6)=12x It doesnt give me 13x

OpenStudy (cwrw238):

you can factor this you need two numbers whose product is + 36 and whose sum is -13

OpenStudy (anonymous):

or you could do the quadratic formula?

OpenStudy (anonymous):

What is that

OpenStudy (cwrw238):

it factors to binomials: (x )(x ) where 2 numbers go into blanks

OpenStudy (anonymous):

\[x=\frac{ -b \pm \sqrt{b^2 -4ac} }{ 2a }\] Where a, b and c are the coefficients in the quadratic (here a=1, b=-13 and c = 36) So if you put these numbers in you will get x=9 and x=4.

OpenStudy (anonymous):

(x-4)(x-9)

OpenStudy (cwrw238):

think of factors of + 36 4 and 9 you need -13 so its -4 and -9 factors are (x - 4)(x - 9)

OpenStudy (anonymous):

then from there you just solve for 0

OpenStudy (anonymous):

so it would be 4 and 9

OpenStudy (anonymous):

yep (:

OpenStudy (cwrw238):

if you have equation x^2 - 13x + 36 = 0 the roots are 4 and 9

OpenStudy (cwrw238):

when you ask to solve it usually means solve an equation so you should have written it as an equation that is x^2 - 13x + 36 = 0